### v1

Abstract
Market definition holds significant importance in antitrust cases, yet achieving consensus on the correct approach remains elusive. As a result, analysts routinely entertain multiple market definitions to ensure the resilience of their conclusions. I propose a simple framework for conducting organized sensitivity analysis with respect to market definition. I model candidate market definitions as partially ordered and use a Hasse diagram, a directed acyclic graph representing a finite partial order, to summarize the sensitivity analysis. I use the Shapley value and the Shapley-Shubik power index to quantify the average marginal contribution of each firm in driving the conclusion. I illustrate the method’s usefulness with an application to the Albertsons/Safeway (2015) merger.
Keywords: Market definition, sensitivity analysis, partial order, Hasse diagram, Shapley value
1 Introduction
Over the course of U.S. antitrust litigation history, the resolution of a greater number of cases has hinged on market definition, more so than any other significant aspect (Baker,, 2007). Courts in the United States and Europe require competition authorities to define “markets” before assessing competitive effects, as highlighted in various legal cases and statutes. For example, the FTC-DOJ Merger Guidelines have used both the level and the change in Herfindahl-Hirschman Index (HHI)—which depend on market definition—to determine whether a proposed merger should be presumed illegal. [^3] In their recently released 2023 Merger Guidelines, the FTC and DOJ changed the HHI thresholds for triggering structural presumption.
Unfortunately, the use of market definition entails two critical complications. First, the term market definition bears much ambiguity, and there is little consensus on its definition (Glasner and Sullivan,, 2020). A growing contingent of critics have argued for scrapping the market definition altogether and focusing solely on competitive effects analysis, which need not require arbitrary market delineation. [^4] The hostility to market delineation grew from the early 1930s with the development of the theory of monopolistic competition, as product differentiation makes defining a relevant market “an inevitably artificial line-drawing exercise” (Werden,, 1992; Farrell and Shapiro,, 2010). Louis Kaplow, in a series of papers (Kaplow,, 2010, 2011, 2012, 2013, 2015, 2021, 2022), argues that market definition serves no role except to produce market shares, which themselves serve as poor measures of market power, and that the logic of market delineation is inherently circular. While acknowledging the limitations to market delineation, proponents argue that the practice of market delineation has practical merits on theoretical and empirical grounds (Coate and Simons,, 2012; Werden,, 2013; Nocke and Schutz,, 2018; Glasner and Sullivan,, 2020; Miller and Sheu,, 2021; Nocke and Whinston,, 2022; Nocke and Schutz,, 2023). However, the Supreme Court recently reaffirmed that the practice will not be abandoned soon: In Ohio v. American Express, 585 US (2018), the Court disagreed with the plaintiffs’ argument that market definition was not required because they had demonstrated adverse competitive effects. [^5] The Court disagreed and explained that “courts usually cannot properly apply the rule of reason without an accurate definition of the relevant market,” thereby ruling that the plaintiffs failed to establish a prima facie case. See the Court opinion at https://www.supremecourt.gov/opinions/17pdf/16-1454_5h26.pdf. The case clarified when market definition is needed but not how it ought to be defined.
Second, a single antitrust case may be subject to multiple market definitions. [^6] Glasner and Sullivan, (2020) criticizes the common expectation of courts and practitioners that every case should involve one single relevant market as the single market fallacy. Multiple theories of harm may be relevant. Moreover, the popular Hypothetical Monopolist Test may produce multiple market definitions (Baker,, 2007; Davis and Garcés,, 2009; Salop and Moresi,, 2009). Perhaps most importantly, plaintiffs and defendants face uncertainty over which market definition will prevail in court. Thus, scrutinizing the implications of alternative market delineation is a norm rather than an exception. Despite the aforementioned debates and the pervasiveness of sensitivity analysis, attempts to establish a systematic approach have been lacking. This paper aims to fill this gap.
In this paper, I propose a simple framework for conducting an organized sensitivity analysis with respect to market definition. [^7] Sensitivity analysis studies the relative importance of different input factors on the model output. Early advocates such as Leamer, (1985) have recognized the need for organized sensitivity analysis, which has led to the development of the literature on semiparametric models that weaken parametric assumptions or set-identified models that give up on point identification. Leamer, (1985) argues that “[a] fragile inference is not worth taking seriously…. Conclusions are judged to be sturdy only if the neighborhood of assumptions is wide enough to be credible and the corresponding interval of inferences is narrow enough to be useful.” In antitrust settings, plaintiffs and defendants have strong incentives to employ extensive sensitivity analysis to present the robustness of their results with respect to market definitions and determine where to allocate their investigative resources. The main contributions are twofold. First, I model the sensitivity analysis problem by leveraging the partial order nature of candidate markets and use Hasse diagram—a directed acyclic graph for representing a finite partial order—to visualize the sensitivity of metrics of interest with respect to market definitions. My diagrammatic approach facilitates a compact summary of multi-dimensional measures from the sensitivity analysis. [^8] Mathematical diagrams are used in various branches of science to summarize information. The diagrammatic approach to causal inference has proven extremely useful; see, e.g., Pearl, (1995). I provide a novel application of Hasse diagrams for sensitivity analysis. Second, I measure the importance of each firm (or a group of products/firms) using the Shapley value and the Shapley-Shubik power index. [^9] In its original form, the Shapley value is a solution concept in cooperative game theory that finds a fair distribution of gains by measuring each player’s contribution. Shapley-Shubik power index is the Shapley value defined on simple games in which the outcome of a coalition is binary (0 or 1). I interpret the removal of firms from an antitrust market as the firms “participating in a coalition” to generate the final evaluation metric of interest (e.g., Herfindahl-Hirschman index or gross upward pricing pressure). Shapley value allows the analyst to understand the importance of each firm in determining the final conclusion.
I illustrate the framework’s usefulness with an application to the Albertsons/Safeway merger consummated in 2015. Evaluating a supermarket merger requires drawing boundaries on which grocery chains should be considered sufficiently close competitors to traditional supermarkets. Many store formats serve grocery products to consumers but operate on heterogeneous business models. I examine the sensitivity of market concentration measures with respect to the inclusion and exclusion of wholesale club, natural/gourmet, and limited assortment stores. I also decompose the contribution of each major chain in triggering the structural presumption defined by the 2023 FTC-DOJ Merger Guidelines.
This paper contributes to the large literature on antitrust market definitions; see Werden, (1992), Baker, (2007), and Davis and Garcés, (2009) for reviews. While much effort has been dedicated to identifying the “right” market definition, the literature has paid less attention to the sensitivity analysis issue. I fill this gap by developing novel quantitative tools for conducting organized sensitivity analysis with respect to market definition. I find a novel application of Shapley value, a well-known cooperative game theory solution concept primarily applied to study voting power. The use of Shapley value for sensitivity analysis has been increasing in the statistical and machine learning literature (Owen,, 2014; Song et al.,, 2016; Plischke et al.,, 2021; Rozemberczki et al.,, 2022). To the best of my knowledge, quantitative sensitivity analysis with respect to market definitions is new.
My framework also complements empirical tools that hinge on market definitions. While I focus on HHI as an evaluation metric throughout the paper for simplicity, my framework extends to various statistics that may be of interest, including concentration ratio, Herfindahl-Hirschman index (Hirschman,, 1964; Nocke and Schutz,, 2018; Nocke and Whinston,, 2022; Nocke and Schutz,, 2023), merger simulation (Nevo, 2000a, ), compensating marginal cost reduction (Werden,, 1996), upward pricing pressure (Farrell and Shapiro,, 2010; Conlon and Mortimer,, 2021; Miller and Sheu,, 2021; Koh,, 2024). [^10] In principle, upward pricing pressure and compensating marginal cost reductions need not rely on market definition (if reliable diversion ratio data are available). On many occasions, however, they do because diversion ratios need to be estimated. First, especially in the merger screening stage, it is common for economists to specify the demand function as logit so that diversion ratios can be approximated with the proportional-to-share formula, which depends directly on market definition. Second, even flexible demand models (e.g., random-coefficients logit models) can be sensitive to market size parameters or the set of products included in consumers’ consideration set, which relate to market definition (Nevo, 2000b, ). More broadly, the framework may be applied to other settings where the analyst wants to report the sensitivity of the main results with respect to a partially ordered discrete set of assumptions. [^11] For example, a Hasse diagram can be useful in summarizing how the size of the partially identified sets with respect to alternative assumptions on players’ information and solution concept in discrete game models; see, e.g., Koh, (2023). In the field of empirical industrial organization, it is also standard to consider multiple market definitions or explore the implication of different nesting structures in nested logit models. Note that this paper makes no attempts to select the right market definition nor respond to the current debate over market definition.
The rest of the paper is organized as follows. Section 2 presents a model of sensitivity analysis and the use of Hasse diagrams. Section 3 illustrates how the Shapley value can measure the importance of each firm in market definition. Section 4 considers an empirical application. Section 5 concludes.
2 A Model of Sensitivity Analysis
Expanding the product and/or geographic markets adds more competitors and reduces each firm’s market share. Many popular statistics for evaluating market power and competitive effects of mergers—e.g., concentration ratio, Herfindahl-Hirschman index (both in absolute level and changes from pre-merger level to post-merger level), compensating marginal cost reduction (Werden,, 1996), and upward pricing pressure (Farrell and Shapiro,, 2010)—are functions of market shares, and expanding market definition generally leads to more conservative (lower) estimates. [^12] Again, this statement applies to upward pricing pressure and compensating marginal cost reductions when demand function is specified as logit or CES. For ease of exposition, I focus on a scenario in which the analyst is interested in examining a model’s sensitivity to the exclusion of potential competitors so that expanding the list of firms to be removed from the market increases the evaluation metric.
Let $\mathcal{N}\equiv\{1,2,...,N\}$ be the set of marginal firms under consideration. Each $\omega\in 2^{\mathcal{N}}$ represents an exclusion set defined as the list of firms to be dropped from the market. Function $f:2^{\mathcal{N}}\to\mathbb{R}$, which I refer to the outcome function, returns the value of an evaluation metric. The analyst’s objective is to study the sensitivity of $f$ with respect to changing $\omega$. I provide a running example below.
Example 1.
Suppose the universe of all firms is $\{A,B,1,2,3\}$. Firms A and B announce a merger. To evaluate its impact on market concentration, the analyst wants to compute HHI but is unsure whether Firms 1, 2, or 3 should enter the relevant market definition. In this case, the set of marginal members under consideration is $\mathcal{N}=\{1,2,3\}$. The function $f(\omega)$ measures the HHI excluding firms in $\omega\in 2^{\{1,2,3\}}$ from $\{A,B,1,2,3\}$. For example, $f(\{1,3\})$ measures the HHI after taking the relevant market to be $\{A,B,2\}$. ∎
2.1 Hasse Diagram Representation of Finite Partial Order
The partial order $(2^{\mathcal{N}},\subseteq)$, where $\subseteq$ represents the inclusion relation, formalizes the idea that a market definition may be broader or narrower than another, but two markets need not be comparable. To graphically represent the partial order, I use a Hasse diagram, a mathematical diagram representing a transitive reduction ordering of a finite partially ordered set.
Definition 1 (Hasse diagram).
[…]
Any finite partial order admits a Hasse diagram representation; however, a Hasse diagram representation for a given finite partial order may not be unique. I provide an example of a Hasse diagram as follows.
Example 2 (Continued).
Figure 1 is a Hasse diagram for $(2^{\mathcal{N}},\subseteq)$ where $\mathcal{N}=\{1,2,3\}$. With nodes defined by the elements of $2^{\mathcal{N}}$ and edges defined by the inclusion relation, the directed acyclic diagram compactly summarizes the hierarchical relationships among the subsets of $\{1,2,3\}$. The bottom node $\emptyset$ represents the broadest market where no firm is excluded. The top node $\{1,2,3\}$ corresponds to the narrowest market in which all firms in $\{1,2,3\}$ are excluded, and only the merging firms $\{A,B\}$ remain. Note that edges are drawn only between nodes that have a covering relationship; for example, there is no direct arrow from $\{1\}$ to $\{1,2,3\}$ since $\{1,2\}$ (or $\{1,3\}$) is positioned between them.
Figure 1: An Example of Hasse Diagram
It is always possible to draw a Hasse diagram with a subset of nodes since any subset of nodes form a partial order, and any finite partial order admits a Hasse diagram representation. Drawing a Hasse diagram with a subset of nodes can be helpful when the power set of $\mathcal{N}$ is too large and when particular subsets of firms are of primary interest.
2.2 Measures of Competitive Effects
It is possible to augment the diagram by layering each node with the value of the outcome function. In a typical application, $f:2^{\mathcal{N}}\to\mathbb{R}$ is monotonic. In Appendix A, I establish the monotonicity of popular statistics (concentration ratios, HHI, upward pricing pressure, and compensating marginal cost reduction) with respect to market definition. [^14] I emphasize again that while upward pricing pressure and compensating marginal cost do not rely on market definition (if reliable diversion ratio data are available), practitioners often rely on logit demand assumptions to approximate diversion ratios using cross-sectional market shares data.
Example 3 (Continued).
Suppose that there are a total of eight firms with sales $x_{A}=15$, $x_{B}=15$, $x_{C}=10$, $x_{D}=10$, $x_{E}=10$, $x_{1}=9$, $x_{2}=6$, $x_{3}=3$. I assume firms $\{A,B,C,D,E\}$ are always included in the market but test the sensitivity of HHI with respect to the exclusion of Firms 1, 2, and 3. Figure 2 summarizes the sensitivity of pre-merger HHI with $\mathcal{N}=\{1,2,3\}$. Each node reports the level of HHI for a given market definition, and each edge reports the marginal change in HHI as the market definition narrows. For example, the HHI associated with the broadest market definition (bottom node) is 1439. However, excluding Firm 1 increases the HHI by 230 to 1669. Excluding Firms 1, 2, and 3 leads to an HHI of 2083. In general, as shown in Appendix A, excluding more firms in the relevant market increases HHI insofar as the excluded firms’ sales are not too large.
Figure 2: Annotated Hasse Diagram
I assume that the HHI threshold for a highly concentrated market is 1,800, as defined in the 2023 FTC-DOJ Merger Guidelines. In Figure 2, the nodes corresponding to high-HHI markets are shaded in red. It shows that excluding Firm 1 is essential for triggering the high-HHI threshold, but either Firm 2 or 3 needs to be excluded from the market as well. ∎
As seen from the example above, Hasse diagrams can organize complex information based on the partial order of market definitions. Although I have assumed a single outcome function $f$, it is possible to augment the diagram with multidimensional information. Yet, the complexity of the diagram quickly rises as the number of firms in $\mathcal{N}$. Specifically, the Hasse diagram for $(2^{\mathcal{N}},\subseteq)$ has $2^{N}$ nodes and $N\cdot 2^{N-1}$ edges. In the following section, I propose using the Shapley value to measure the average contribution of each firm.
3 Sensitivity Analysis with Shapley Value
Suppose that the outcome function $f$ is monotonic such that the broadest market definition generates the most conservative (lowest) estimate of market power and vice versa. Then how much does each firm contribute in driving the output from $f(\emptyset)$ to $f(\mathcal{N})$? I show that the Shapley value provides a simple answer to this question.
3.1 Shapley Value for Sensitivity Analysis
I begin by reviewing the definition of coalitional games and Shapley values as follows.
Definition 2 (Shapley value).
A pair $\langle\mathcal{N},v\rangle$ is a coalitional game if $\mathcal{N}$ is a finite set of players and $v:2^{\mathcal{N}}\to\mathbb{R}$ is a characteristic function such that $v(\emptyset)=0$. The Shapley value for each player $i$ in a coalitional game $\langle\mathcal{N},v\rangle$ is
$\varphi_{i}=\sum_{S\subseteq N\backslash\{i\}}\frac{|S|!(N-|S|-1)!}{N!}(v(S\cup\{i\})-v(S)).$ (1)
Shapley value (Shapley,, 1953) is a single-valued solution concept for coalition games that proposes how to distribute the total gains to the players, assuming that they all collaborate. [^15] Shapley value is defined by an axiomatic approach: it is the unique solution concept that satisfies the efficiency, symmetry, null player, and additive properties. See Maschler et al., (2020) for a textbook treatment of the topic. In Definition 2, the function $v(S)$ describes the value of coalition $S\in 2^{\mathcal{N}}$. For each $S\in 2^{\mathcal{N}}$ such that $i\notin S$, $v(S\cup\{i\})-v(S)$ is the marginal surplus that player $i$ brings to the table. The Shapley value of $i$ is the average marginal contribution over all possible $S$ that permits such calculation. Thus, (1) measures the average marginal contribution of each player over the possible different permutations in which the coalition can be formed. A key property of Shapley value is that
$\sum_{i=1}^{N}\varphi_{i}=v(\mathcal{N}),$ (2)
i.e., the sum of Shapley values is equal to the total value of the grand coalition. Shapley value measures how much the players can expect from a fair distribution of total surplus generated by the grand coalition.
Shapley value has a natural application to sensitivity analysis because it decomposes the contribution of each member in generating the final output. I interpret the removal of firms in the relevant market as the firms forming a coalition. For example, if firms 1 and 2 are removed from the market definition, it is as if they formed a coalition to exclude themselves from the market. Since the set of firms $\mathcal{N}=\{1,...,N\}$ is given, the description of a coalitional game is completed by defining a characteristic function. For each subset of firms $S\in 2^{\mathcal{N}}$, define the characteristic function as
$v(S)\equiv f(S)-f(\emptyset).$ (3)
The characteristic function (3) is simply an outcome function normalized so that $v(\emptyset)=0$. The Shapley value measures the importance of each firm in increasing the outcome measure by $v(\mathcal{N})$ when going from the broadest market definition to the narrowest market definition.
Example 4 (Continued).
Recall from Figure 2 that $f(\emptyset)=1439$ and $f(\{1,2,3\})=2083$. Thus, excluding firms $\{1,2,3\}$ from the market increases the level of HHI by $v(\{1,2,3\})\equiv f(\{1,2,3\})-f(\emptyset)=2083-1439=644$. Table 1 reports the Shapley values of each firm. [^16] I use R package CoopGame (Staudacher and Anwander,, 2019) to compute the Shapley values. Note that the Shapley values sum to the value of grand coalition 644. Column “SV Share” says that Firms 1, 2, and 3 account for 44%, 35%, and 21% of the total increment. The share of average contribution measured by Shapley value differs from the sales shares, which are 50%, 33%, and 17% for Firms 1, 2, and 3, respectively. Using the sales shares to explain the importance of each firm in the HHI calculation overstates Firm 1’s contribution and understates Firm 3’s contribution.
Table 1: Shapley Value and Shapley Shubik Power Index
Firm Sales Sales Share SV SV Share SSPI
1 9 50% 282 44% 0.666
2 6 33% 228 35% 0.166
3 3 17% 134 21% 0.166
Total 60 100% 644 100% 1.0
Table 1 is significant because it breaks down each firm’s contribution to the final outcome measure of interest. In this example, while sales shares and Shapley values are roughly proportional, using Shapley values offers a distinct advantage. The analyst derives the Shapley values directly from the outcome values, which depend nonlinearly on shares, whereas sales shares merely preserve rankings. In the following, I demonstrate that the utility of Shapley values is even greater when the outcome measure represents a binary decision rule, which is ultimately what matters.
3.2 Shapley-Shubik Power Index
I show that Shapley-Shubik Power Index (Shapley and Shubik,, 1954) can quantify the relative importance of each member in triggering a binary decision such as the structural presumption for merger screening.
Definition 3.
A coalitional game in which the worth of each coalition is 0 or 1 is called a simple game, i.e., the outcome function is $f:2^{\mathcal{N}}\to\{0,1\}$. When applied to simple games, the Shapley value is known as the Shapley-Shubik power index.
Intuitively, the Shapley-Shubik power index of a player is the probability that the player will be a pivot player; in this sense, the index measures the power of each player in turning the outcome from 0 (“loss”) to 1 (“win”). Shapley-Shubik power index is widely used in political science as a measure of power distribution in committees.
Many antitrust applications can be cast as simple games because certain binary decisions are triggered depending on how the relevant market is defined. For example, the 2023 FTC-DOJ Merger Guidelines presume that a merger substantially lessens competition or tends to create a monopoly if either (i) the post-merger HHI is greater than 1,800, and the merger increases the HHI by more than 100, or (ii) the merged firm’s market share is greater than 30%, and the merger increases the HHI by more than 100. The analyst can encode the outcome as 1 if a given market definition triggers the structural presumption and 0 otherwise. Removing more competitors will likely increase the chance of triggering the structural presumption. The Shapley-Shubik power index measures the importance of each member in triggering the presumption.
Example 5 (Continued).
Suppose the analyst wants to determine whether the pre-merger market is highly concentrated. A market is deemed highly concentrated if the HHI is higher than 1800. Recall that $v(S)$ measures the pre-merger HHI associated with coalition $S$. The decision rule induces a simple game with characteristic function
$v^{*}(S)=\begin{cases}1&\quad\text{if}\quad v(S)\geq 1800,\\ 0&\quad\text{if}\quad v(S)<1800.\end{cases}$ (4)
Column “SSPI” of Table 1 reports the Shapley-Shubik power index derived from $(\mathcal{N},v^{*})$. First, it gives Firm 1 a weight of 67%, far higher than the weights of the other two firms. Second, surprisingly, Firms 2 and 3 have the same power index values, indicating that the two firms have equal importance in triggering the decision even though one firm has twice the market share as the other. Thus, the example illustrates why it is incorrect to assume that the importance of each firm in triggering a decision is strictly proportional to market shares. ∎
4 Empirical Application to Albertsons/Safeway (2015)
In this section, I apply the proposed framework to the Albertsons/Safeway merger. I analyze the sensitivity of the HHI calculation and structural presumption with respect to the exclusion of club, natural, and limited assortment stores, which may or may not be considered close competitors to traditional supermarkets. [^17] I make no claims on which formats should be included in the relevant market definition when evaluating supermarket mergers. I consider both state-level and local-level analyses (see below for the definition of local markets).
4.1 Background
In March 2014, the Albertsons supermarket chain entered into an agreement to purchase Safeway for \$9 billion. Albertsons operated 1,075 stores in 28 states under the banners Albertsons, United, Amigos, and Market Street, among others; Safeway owned 1,332 stores in 18 stores under the banners Safeway, Vons, Pavilions, Tom Thumb, and Randall’s, among others. If consummated, the merger was to create the second-largest traditional grocery chain by store count and sales in the US. In late 2014, the FTC settled with the parties with a mandate to divest stores and production facilities.
The FTC defined the relevant product market as supermarkets within “hypermarkets.” [^18] See FTC’s Analysis of Agreement Containing Consent Order to Aid Public Comment available at https://www.ftc.gov/system/files/documents/cases/150127cereberusfrn.pdf. Supermarkets refer to “traditional full-line retail grocery stores that sell, on a large-scale basis, food and non-food products that customers regularly consume at home—including, but not limited to, fresh meat, dairy products, frozen foods, beverages, bakery goods, dry groceries, detergents, and health and beauty products.” Hypermarkets include chains such as Walmart Supercenters that sell an array of products not found in traditional supermarkets but also offer goods and services available at conventional supermarkets. The relevant geographic markets were defined as areas that range from a two- to ten-mile radius around each of the defendants’ stores, where the relevant radius varies depending on factors such as population density, traffic patterns, and unique characteristics of each market.
Historically, defining the relevant antitrust market for supermarkets has been challenging for several reasons. Supermarkets offer a wide range of products beyond groceries, including household items, pharmacy goods, and electronics. Consumers also shop at multiple types of stores for groceries, including wholesale clubs (e.g., Costco, Sam’s Club), limited assortment stores (e.g., Aldi, Lidl), natural/gourmet stores (e.g., Whole Foods, Sprouts), and ethnic stores (e.g., H Mart, Fiesta Mart). Traditionally, the FTC defined the relevant market as “full-line grocery stores” that provide “one-stop shopping” that enables consumers to “shop in a single store for all of their food and grocery needs.” [^19] See, for example, Albertsons/Safeway, Ahold/Delhaize, and Whole Foods/Wild Oats mergers. The debate over market definition was particularly intense during Whole Foods/Wild Oats: While the FTC contended that the relevant market should be defined as “premium, natural, and organic supermarkets,” the merging parties asserted that the relevant product market should include all supermarkets.
4.2 Data
I illustrate how the proposed approaches can facilitate sensitivity analysis with respect to market definition. I study grocery chains’ market shares in the State of Washington, where Albertsons and Safeway had a significant overlap in their footprint. I use the 2009 cross-section of the AC Nielsen’s (currently known as The Nielsen Company) Trade Dimensions data, which reports retail stores’ locations, revenues, and characteristics. [^20] Year 2009 is the closest year before 2014 for which I have access to the Trade Dimensions data. To mimic the Albertsons/Safeway merger investigation that started in 2014, I encode Albertsons stores in Washington State as owned by Albertsons; SuperValu Inc. had ownership over the Albertsons banner in the region but sold it to Cerberus Capital Management (which owned the Albertsons company) in 2013. The Trade Dimensions dataset categorizes each grocery store into one of multiple formats. I categorize stores into “supermarket,” “supercenter,” “club,” “natural,” and “limited.” [^21] Specifically, I categorize a store as a “supermarket” if it is categorized as a supermarket or a superette (small supermarkets) in the Trade Dimensions data; “supercenter” if categorized as a supercenter; “club” if categorized as a wholesale club or a warehouse; “natural” if categorized as a natural/gourmet store; “limited” if categorized as a limited assortment tore. I include all grocery stores as potential competitors but drop military commissary stores.
Table 2: Summary Statistics
Variable Mean Median Min Max
Revenue (\$mil) $17.23$ $13.04$ $0.00$ $95.16$
Supermarket $0.74$ $1.00$ $0.00$ $1.00$
Supercenter $0.10$ $0.00$ $0.00$ $1.00$
Club $0.07$ $0.00$ $0.00$ $1.00$
Natural/Gourmet $0.06$ $0.00$ $0.00$ $1.00$
Limited Assortment $0.04$ $0.00$ $0.00$ $1.00$
  • Notes: The sample consists of all grocery stores in Washington State (excluding military commissary stores).
Table 2 reports the basic summary statistics. In 2009, an average grocery store earned \$17.23 million in revenue, but the revenues could vary substantially across stores. The dominant formats (in store counts) were traditional supermarket and supercenter, which accounted for 74% and 10% of all stores, respectively. Club, natural, and limited assortment stores accounted for less than 20% of all Washington State stores.
Table 3: Top 11 grocery chains in Washington State
Firm Primary Format Revenue (\$mil) Market Share Store Count
Kroger Supermarket $3239.37$ $0.217$ $128$
Safeway Supermarket $3058.17$ $0.205$ $168$
Costco Club $1862.80$ $0.125$ $29$
Walmart Supercenter $1805.44$ $0.121$ $34$
Albertsons Supermarket $1062.41$ $0.071$ $75$
Winco Supermarket $508.39$ $0.034$ $12$
Haggen Supermarket $449.73$ $0.030$ $29$
Trader Joes Natural $241.16$ $0.016$ $15$
Whole Foods Natural $181.19$ $0.012$ $5$
Grocery Outlet Limited $170.71$ $0.011$ $30$
Rosauers Supermarket $140.26$ $0.009$ $14$
Table 3 reports the revenue from grocery sales, market share, and the number of stores for the top 11 grocery chains in Washington State. The market shares in the table are calculated using all grocery stores in the sample. The top five grocery chains account for almost three-quarters of the total revenue in the sample. Costco and Walmart combined had a significant presence with approximately 25% market share in revenue. Trader Joe’s and Whole Foods were the largest natural grocery retail chains. Grocery Outlet was the largest limited assortment chain.
4.3 State-Level Analysis
As a first exercise, I ignore the spatial nature of grocery competition and calculate chains’ market shares at the state level. The primary outcome measure of interest is whether the merger triggers structural presumptions. In this exercise, following the 2023 FTC-DOJ Merger Guidelines, I assume that a merger is presumed illegal if the post-merger HHI is above 1,800 and the HHI increases by more than 100.222222The 2023 Merger Guidelines presumes that a proposed merger is illegal if either (i) post-merger HHI is above 1,800 and the HHI increases by more than 100, or (ii) the merged firm’s market share is above 30% and the HHI increases by more than 100. For simplicity, I only apply the first criterion. I examine the sensitivity of the outcomes with respect to grocery store formats. I always include supermarkets and supercenters in the market definition. I test how the outcomes change when I exclude club stores, limited assortment stores, or natural stores. That is, I set $\mathcal{N}=\{\text{Club},\text{Natural},\text{Limited}\}$.
Figure 3 illustrates how the post-merger HHI and $\Delta$HHI (defined as the difference between post-merger HHI and pre-merger HHI) vary as I exclude more formats from the market. Each pair in the nodes represents the values of post-merger HHI and $\Delta$HHI under the assumed market definition. Nodes shaded in red correspond to market definitions that trigger the structural presumption. The values in the edges represent the changes in post-merger HHI and $\Delta$HHI when shrinking the market definition by excluding certain formats.
Figure 3: Post-Merger HHI and $\Delta$HHI
When I include all grocery formats (supermarkets, supercenters, clubs, limited assortment stores, and natural stores), the Albertsons/Safeway merger produces the post-merger HHI of 1,576 and $\Delta$HHI of 293 in Washington State. [^23] Post-merger HHI of 1,576 falls short of being labeled as “highly concentrated” in the 2023 Merger Guidelines. However, the 1982 version classified HHI between 1,000 and 1,800 as “moderately concentrated” until 2010. Excluding club stores triggers the structural presumption. However, excluding natural or limited assortment stores from the market definition does not affect the structural presumption, as doing so only slightly increases the post-merger HHI. Removing all club, natural, and limited stores raises the post-merger HHI and $\Delta$HHI to 2,152 and 456, respectively.
Table 4 reports the post-merger HHI Shapley value of each format. Out of the total increment of 576 ($=2152-1576$) in the post-merger HHI resulting from removing the three formats, club, natural, and limited assortment formats account for 64.0%, 26.8%, and 9.2%, respectively. Thus, club format alone accounts for more than half of the increase in post-merger HHI, even though the number of club stores is small relative to other formats. The SSPI column, which reports the Shapley-Shubik power index of each format, reveals that club format is a “dictator”: triggering the structural presumption only depends on whether club stores are included; natural and limited assortment formats have no role.
Table 4: Shapley Values for Post-merger HHI and Structural Presumption
Format SV SV Share SSPI
Club 386.71 0.640 1
Natural 154.23 0.268 0
Limited 52.86 0.092 0
Total 575.80 1.0 1.0
From the state-level analysis, the analyst can infer that only club stores merit attention regarding whether they are in the relevant antitrust market; investigative resources should be directed to examining whether consumers find merging parties’ stores and club stores as close substitutes. In Appendix B, I apply the Shapley-Shubik power index to analyze the importance of each competitor firm in triggering the structural presumption. This exercise illustrates that the Shapley values and the Shapley-Shubik power index can be calculated with many players.
4.4 Local Market Analysis
Consumers mostly shop at supermarkets around residences or workplaces as traveling is costly. As such, competition in the supermarket industry is local in nature. The FTC defined the relevant geographic markets as areas that range from a two- to ten-mile radius around each supermarket. As no further details are provided in the FTC’s publicly released complaints, I simply define local “circle” markets as all grocery stores within a 5-mile radius around each of the defendant’s supermarkets. [^24] In other words, each Albertsons/Safeway store forms a 5-mile circle market that can be interpreted as a local geographic market. Defining local markets as a collection of stores around each “center store” is useful because the analyst does not need to define arbitrary geographic boundaries. If a local geographic market, however defined, does not have a high level of market concentration or significantly raise concentration, circle markets defined around center stores will likely not raise a presumption of illegality as well. In practice, antitrust authorities will consider specific geographic characteristics (e.g., rivers or highways) to determine the relevant geographic market and adjust the radius according to the urbanicity of local markets. I calculate the post-merger HHI and $\Delta$HHI in each local market.
Figure 4: Number of Presumptive Local Markets
Figure 4 shows the number of presumptive local markets. There are 143 presumptive markets when all store formats are included in the market definition. However, the number increases to 188 when club, natural, and limited assortment stores are excluded from the market definition (only supermarkets and supercenters are considered), indicating 45 markets are sensitive to market definition. Removing club stores from the market definition significantly raises the number of presumptive markets, indicating that club stores deserve major attention.
Table 5 reports the frequency of SSPI structure for the 45 sensitive markets. For example, out of the 45 sensitive markets, 18 markets have $(\text{SSPI}_{\text{Club}},\text{SSPI}_{\text{Natural}},\text{SSPI}_{\text{Limited}})=(1,0,0)$, meaning whether the structural presumption is met only depends on the exclusion club stores (club format is the “dictator” in these local markets). The distribution of SSPIs across the local markets indicates that the club format tends to have a high power index. Figure 5 visualizes Table 5 (random noises were added to the points to enhance visibility). It also shows that clubs and natural stores have a high power index. In contrast, limited assortment stores are less important.
Table 5: Local Market SSPI
SSPI Structure
Club Natural Limited Count
1.000 0.000 0.000 18
0.667 0.167 0.167 4
0.500 0.500 0.000 14
0.500 0.000 0.500 2
0.333 0.333 0.333 3
0.167 0.667 0.167 1
0.000 1.000 0.000 3
Figure 5: Shapley-Shubik Power Index of Local Markets
The Shapley-Shubik power index shows that whether the structural presumption is met largely depends on whether club stores can be removed from the market definition. Exclusion of natural stores can also have a high impact, but limited assortment stores tend to matter little. In sum, the sensitivity analysis reveals which markets are sensitive to market definition, which formats are more important than others, and how investigative resources should be allocated to defend the conclusions on merger illegality.
5 Conclusion
Attempts to pinpoint the right market definition often prove fruitless. The FTC-DOJ Merger Guidelines acknowledge that a single case may be subject to multiple market definitions. Even in the most ideal case, the adversarial nature of litigation requires the plaintiffs, defendants, and court to scrutinize the sensitivity of the conclusion with respect to market definition.
Contrary to existing works that attempt to find the best approach for defining an antitrust market, I take the possibility of entertaining multiple market definitions as given and propose a simple framework for conducting organized sensitivity analysis. I show that Hasse diagrams can be useful for summarizing the sensitivity of model outcomes with respect to market definition. Moreover, the Shapley value and the Shapley-Shubik power index are simple tools for quantifying the impact of including or excluding firms from the relevant market.
Although I have focused on a particular application, the proposed idea can be applied more broadly. For example, estimating economic margins is also an area of contention as margins are key inputs to competitive effects analysis; the plaintiffs and defendants may disagree on the cost categorization, which naturally leads to the need for sensitivity analysis. More generally, a researcher may use the proposed tools to evaluate the sensitivity of policy recommendations with respect to a set of discrete assumptions. The framework also relates to the partial identification literature, which has tight connections to sensitivity analysis. I leave further development and applications to future work.
Appendix A Monotonicity of Popular Evaluation Metrics
Removing products or firms from the market narrows the market definition and increases the market shares of each remaining firm. I show that popular measures of market power and competitive effects that depend on market shares have the monotonicity property under the logit demand assumption. To set the stage, let $m$ denote a market that consists of a set of firms. If $m^{1}\supseteq m^{0}$, then $m^{1}$ is a broader market. Let $x_{j}$ represent firm $j$’s sales. The share of firm $j$ is $s_{j,m}\equiv x_{j}/\sum_{l\in m}x_{l}$.
A.1 Concentration Ratios
The concentration ratio is calculated as the sum of the market share percentage held by the largest specified number of firms in an industry. Define the concentration ratio of a group $g\subseteq m$ as
$f(g;m)=\sum_{j\in g}s_{j,m}.$
If $m^{1}\supseteq m^{0}$, then $s_{j,m^{0}}\geq s_{j,m^{1}}$ for all $j\in m^{0}$. It follows that $f(g;m^{0})\geq f(g;m^{1})$, i.e., defining a smaller market leads to a higher concentration ratio statistic.
A.2 Upward Pricing Pressure
In a single-product firm setting, Farrell and Shapiro, (2010) defines the upward pricing pressure of firm $j$ merging with firm $k$ is defined as
[…]
where $(p_{k}-c_{k})$ is the pre-merger price-cost margin of firm $k$ and $D_{j\to k}$ is the diversion ratio from $j$ to $k$. Under the logit demand assumption, diversion ratios are proportional to share:
$D_{j\to k}=\frac{s_{k}}{1-s_{j}}.$
The numerator is increasing in $s_{k}$ and the denominator is decreasing in $s_{j}$. Thus, defining a smaller market increases the share of firms $j$ and $k$, leading to higher diversion ratio estimates. It follows that defining a smaller market leads to a higher estimate of upward pricing pressure.
A.3 Compensating Marginal Cost Reductions
Werden, (1996) shows that the compensating marginal cost reduction for firm $j$ is
$\ddot{c}_{j}=\frac{m_{j}D_{j\to k}D_{j\to k}+m_{k}D_{k\to j}p_{k}/p_{j}}{(1-m_{j})(1-D_{j\to k}D_{k\to j})},$
where $m_{l}=(p_{l}-c_{l})/p_{l}$ is the pre-merger margin. Since the numerator is increasing in diversion ratios and the denominator is decreasing in diversion ratios, it follows that the compensating marginal cost reductions increases with smaller market definition under the logit assumption.
A.4 Herfindahl-Hirschman Index
The HHI of market $m$ is defined as
$H_{m}=\sum_{j\in m}s_{j,m}^{2}.$ (5)
For a given market definition $m$ that includes the merging firms $a$ and $b$, the increase in HHI from the merger is
$\Delta H_{m}\equiv H^{\text{post}}_{m}-H^{\text{pre}}_{m}=(s_{a,m}+s_{b,m})^{2}-(s_{a,m}^{2}+s_{b,m}^{2})=2s_{a,m}s_{b,m}.$ (6)
From (6), it is apparent that defining a smaller market increases $\Delta H_{m}$.
However, establishing the monotonicity of HHI (in level) is more subtle because HHI depends on the share of all firms in the given market. It is well-known that HHI is decreasing in the number of firms when they are symmetric (Tirole,, 1988). However, HHI need not be decreasing in the number of firms when firms are asymmetric because including a large firm may result in a higher HHI. [^25] To see this, consider the following example. Suppose that $m=\{1,2,3\}$, $x_{a}=1$, $x_{b}=1$, and $x_{c}=98$. Consider $m^{\prime}=\{1,2\}$ obtained by removing the largest firm $c$. Then $f(m)=9,606$ but $f(m^{\prime})=5,000$, so a narrower market decreases HHI. More generally, driving the sale of the largest firm to infinity makes the HHI converge to the level of a monopoly market. Nevertheless, defining a narrower market leads to a higher HHI insofar as the volume of the marginal firm is sufficiently small.
Consider
$H_{m}=\frac{x_{1}^{2}+...+x_{J}^{2}}{(x_{1}+...+x_{J})^{2}}.$
Suppose that $x_{J}$ is the sale of the marginal firm. I consider how $H_{m}$ behaves when $x_{J}\approx 0$, which mimics leaving $x_{J}$ out of the market but including it in the relevant market. Taking partial derivative with respect to $x_{J}$ yields
$\frac{\partial H_{m}}{\partial x_{J}}=2x_{J}(x_{1}+...x_{J})^{-2}+(x_{1}^{2}+...+x_{J}^{2})(-2)(x_{1}+...+x_{J})^{-3}.$
The above is nonpositive if and only if
$x_{J}(x_{1}+...x_{J})\leq(x_{1}^{2}+...+x_{J}^{2}).$
Thus, if $x_{J}\to 0$, the left-hand side of the inequality approaches zero, whereas the right-hand side remains strictly positive. This shows that $\partial H_{m}/\partial x_{J}$ when $x_{j}\approx 0$, establishing that HHI decreases as the relevant market includes smaller players at the margin.
Appendix B Firm-Level Analysis
Again, using the TDLinx data from 2009, I also examine the impact of the top 9 chains in Washington State in triggering the structural presumption (holding Albertsons and Safeway’s presence fixed). I assume that a market definition triggers a structural presumption if the post-merger HHI is above 1,800 and the change in HHI is above 100. Table 6 reports the Shapley-Shubik power index.
Table 6: Firm-level Shapley-Shubik Power Index
Firm SSPI
Kroger 0.313
Costco 0.313
Walmart 0.313
Winco 0.018
Trader Joe’s 0.011
The Markets 0.011
Whole Foods 0.011
Grocery Outlet 0.011
Natural Grocers 0.000
Total 1.0
Walmart, Kroger, and Costco account for approximately 94%. Surprisingly, Walmart, Kroger, and Costco have the same influence in triggering the structural presumption. Thus, the Shapley value reveals that the importance of a firm in triggering the structural presumption is not strictly proportional to the total sales.